OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 2x + 4y = 12.

OpenStudy (krabsen):

What is y?

OpenStudy (anonymous):

idk

OpenStudy (krabsen):

4y=-2x+12 does that help

OpenStudy (anonymous):

slope = 2; y intercept at (0, 12)

OpenStudy (krabsen):

no, still need to divide by 4 then you will have it :)

OpenStudy (anonymous):

slope =1/2 ; y intercept at (0, 4)

OpenStudy (krabsen):

it is -1/2

OpenStudy (anonymous):

slope = -1/2; y intercept at (0, 3)

OpenStudy (krabsen):

so slope =-1/2 and intercept at (0,4)

OpenStudy (krabsen):

the last i wrote is the anwser :)

OpenStudy (krabsen):

ups sorry you are right :) my bad